OpenStudy (faf3):

I need to create a problem where someone is pole vaulting and one of the first things I'm trying to calculate is projectile motion. please help! i already made up a problem but I'm not sure how to do projectile motion

OpenStudy (mrdoe):

are you familiar with vectors?

OpenStudy (faf3):

yes

OpenStudy (mrdoe):

okay, do you have a specific goal in mind like maximizing the distance or height of the jump?

OpenStudy (faf3):

this is what i have so far, tell me if its ok but I'm open to changing it too: the pole vaulter runs 100m as fast as she can in 1o seconds. this gives her a velocity of 10m/s. The pole vaulter weighs 70kg. The bar is 5 m high. The pole vaulter jumps at an angle of 74 degrees. How long is she in the air? How far foes she travel in the air? this is just what i came up with

OpenStudy (mrdoe):

okay, first, only the x-component of her jump will be 10 m/s, if she needs to clear a 5m pole a more relevant question would be with what force does she need to jump with (in the y-direction)

OpenStudy (mrdoe):

the x-distance question and the time are still good questions

OpenStudy (faf3):

i tried to figure out vix and viy. but it wasn't helping me figure out what i needed i think

OpenStudy (mrdoe):

her running is only contributing to the Vx component, the Vy will come only from her jump

OpenStudy (faf3):

for initial Vx i did 10cos74= 2.76 m/s for intial Vy i did 10sin74= 9.61 m/s

OpenStudy (faf3):

but then thats not for the jump I'm guessing its just for the run to the jump right?

OpenStudy (mrdoe):

if she is running parallel to the ground, than the angle is 0

OpenStudy (mrdoe):

so Vx= 10 m/s and Vy= 10 m/s*sin(0) = 0

OpenStudy (faf3):

is the 0 just for the run before she jumps?

OpenStudy (mrdoe):

asking what her Vy would have to be to clear the 5m pole would be a good question too

OpenStudy (mrdoe):

right, the running is her x-component, the jump is hey y

OpenStudy (faf3):

yes i wanted to figure out what it would take for her to clear the pole too. how would i do that?

OpenStudy (mrdoe):

use the kinematic equations , or use conservation of energy which is generally easier, you can calculate how much energy it would take to reach 5m given her mass and work backwards from there

OpenStudy (faf3):

`these are the equations i have. Vf= Vi + gt change in y= Vit + .5gt^2 change in x= Vxt Vfy^2= Viy^2 + 2g(change in y)

OpenStudy (faf3):

i need help solving it. i tried but i couldn't do it

OpenStudy (mrdoe):

okay, try using energy: her kinetic energy (from her y-velocity) must be equal to her gravitational potential energy at 5m of altitude. Kinetic Energy = (1/2)*mv^2 Gravitational Potential energy (close to earth's surface) = mgh then equate the gravitational potential energy to her kinetic energy, the velocity you are solving for is her Vy component

OpenStudy (mrdoe):

i have to go to class but ill check back later so let me know if your still having trouble

OpenStudy (faf3):

ok thanks! i will still need help with solving the projectile motion of it. because i have to calculate stuff using a minimum of 2 concepts, so i can do energy, projectile motion and a few other things

OpenStudy (mrdoe):

well here you are using potential and kinetic energy as well as conservation of energy, so that 3 :P

OpenStudy (faf3):

energy only counts as one concept for my teacher unfortunately :(

OpenStudy (faf3):

im also going to do the acceleration the vaulter has when sinking in to the cushion, but i also need help.